OpenStudy (diamondanon2):

Please show all work f(x)=x^2+2x+1 find f(x+h) and simplify

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (diamondanon2):

I'm thinking something around x^2+2x+1+h

OpenStudy (faiqraees):

f(x+h) means the output you will get if you input (x+h) in the function So just replace the x's in the function with (x+h)

OpenStudy (skullpatrol):

Hint: replace ALL the x's with x + h

OpenStudy (skullpatrol):

Wherever you see an x put a x+h.

OpenStudy (skullpatrol):

You wrote f(x) + h

OpenStudy (diamondanon2):

so now I have x+h^2+2(x+h) + 1

OpenStudy (skullpatrol):

The x^2 should be (x+h)^2

OpenStudy (faiqraees):

Apply algebraic identity on the first term and expand the second term \[\large\rm (x+h)^2+2(x+h)+1 \]

jhonyy9 (jhonyy9):

or you can make it using that f(x) = x^2 +2x +1 = (x+1)^2 and now just substitute the (x+h) in place of x inside these parentheses and square this so f(x+h) = ((x+h) +1)^2

jhonyy9 (jhonyy9):

@skullpatrol do you agree this please ?

OpenStudy (skullpatrol):

Your method is valid @jhony9 but not "easy" to understand for a beginner :-)

OpenStudy (faiqraees):

@jhonyy9 Your representation is accurate. But still the question requires simplifying so just by simplifying your expression will give her the answer

jhonyy9 (jhonyy9):

thank you Faiq

OpenStudy (diamondanon2):

I kinda get it

OpenStudy (diamondanon2):

so now I have to simplify right?

OpenStudy (faiqraees):

Yes

jhonyy9 (jhonyy9):

ty. skull to you too

OpenStudy (skullpatrol):

@DiamondAnon2 be sure to use lots of parentheses (x+h) is treated as one thing in f(x+h)

OpenStudy (diamondanon2):

okay! thanks all of you :)

OpenStudy (skullpatrol):

Thanks for trying to learn :-)